OpenStudy (lncognlto):

How do I find two consecutive integers between which the root of sin x = 1 - x lies?

OpenStudy (lncognlto):

@Princer_Jones @iGreen

OpenStudy (anonymous):

\[\sqrt{\sin x}=1-x\]like this?

OpenStudy (lncognlto):

Not quite; it means where sin x = 1 - x crosses the x axis.

OpenStudy (lncognlto):

I realise I can just substitute numbers in and find the two between which the sign of the function changes, but it there another way to do it/

OpenStudy (lncognlto):

*?

OpenStudy (phi):

graph y= sin x and y = -x + 1 and see where they are equal (intersect)

OpenStudy (phi):

when graphing sin x , x is in radians

OpenStudy (anonymous):

pi/6 is close to 1/2 and sin of pi/6 is 1/2 so I'd say between 0 and 1

OpenStudy (lncognlto):

Okay, I think I got it. I was wondering I needed to take the derivative or something, but it doesn't seem like it. I also get that the root is between 0 and 1. Thanks!